Review of Chris’s lecture

K. Subramani LDCSEE, West Virginia University, Morgantown, WV

1 Instructions

(i) You are requested to assign points to the various categories identified below.

(ii) The point range for a category is 1-10, with 1 indicating failure to demonstrate passable proficiency and 10 indicating excellence.

(iii) Please provide detailed feedback wherever possible.

2 Review

1. Clarity of Speech

2. Clarity of Content

3. Organization

4. Presentation style 5. Information content

6. Relevance to subject of seminar

7. Additional Feedback - Please be as detailed as possible.